# Efficient Dual-domain Image Dehazing with Haze Prior Perception

### v2

@@ IV-B2 Qualitative Comparisons
Fig. 6 and Fig. 7 show real-world dehazing results on the Dense-Haze [49] and NH-HAZE [50] datasets. In Fig. 6, two dense haze scenes are compared. In the first, our method achieves slightly lower PSNR than Dehamer [34], but surpasses it by 0.183 in SSIM, offering clearer textures and reduced color distortion. In the second scene, with severe haze and significant detail loss, our method restores the overall structure and preserves more details than Dehamer [34].
@@
Fig. 7 compares two non-homogeneous haze scenes. Apart from a slightly lower PSNR in the second scene compared to FocalNet [38], our method outperforms others in all metrics. In the first scene, it minimizes color distortion and avoids heavy artifacts from other methods. In the second scene, it removes haze most effectively, revealing clearer details, though it inadvertently eliminates some snow traces.
@@
Overall, our method achieves the highest SSIM across all real-world cases, demonstrating strong global structure recovery even under severe haze-induced detail degradation.
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
9.79 dB 32.69 dB 34.27 dB 32.18 dB 33.67 dB 32.58 dB 35.20 dB PSNR
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
11.54 dB 34.49 dB 30.01 dB 34.38 dB 40.42 dB 40.29 dB 41.06 dB PSNR
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
22.04 dB 31.72 dB 29.04 dB 28.88 dB 35.11 dB 35.82 dB 39.72 dB PSNR
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
12.35 dB 26.62 dB 28.41 dB 26.86 dB 30.09 dB 31.83 dB 33.39 dB PSNR
(a) Hazy image (b) FFA-Net [25] (c) MAXIM-2S [43] (d) Dehamer [34] (e) FocalNet [38] (f) OKNet [45] (g) Ours (h) GT
Figure 5: Visual comparisons on synthetic hazy images from the SOTS dataset. Key regions highlighted by red boxes are enlarged in the lower-left corner for clearer comparison.
TABLE II: Ablation study for three core modules.
@@
IV-C Ablation Studies
TABLE III: Ablation study for key components in HAFM.
@@ IV-C Ablation Studies
TABLE IV: Ablation study for key components in MGAM.
@@
IV-C1 Effectiveness of Core Modules
To thoroughly evaluate the effectiveness and complementarity of the three core modules, we design four variants: MGAM, HAFM (without feedback correction), HAFM+PCGB (with full feedback-guided correction), and MGAM+HAFM. Table II summarizes the performance of each variant.
As shown in Table II, MGAM improves PSNR by 0.75 dB over HAFM while maintaining the same SSIM of 0.994, highlighting its advantage in enhancing local details. Introducing PCGB into HAFM yields an additional 0.70 dB gain, emphasizing its role in enhancing global degradation modeling via adaptive prior correction. Combining MGAM and HAFM results in a 1.80 dB PSNR improvement over MGAM alone and 2.55 dB over HAFM, benefiting from the synergy between multi-scale spatial feature extraction and dual-domain modulation. Further adding PCGB achieves the best overall performance, confirming that the three modules complement each other and jointly contribute to both local and global restoration with minimal additional computational cost.
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
(a) Hazy image (b) Dehamer [34] (c) Ours (d) GT
Figure 6: Visual comparisons on real-world hazy images from the Dense-Haze dataset. Key regions marked with red boxes are enlarged in left-to-right order and arranged vertically on the right.
(a) Hazy image (b) Dehamer [34] (c) FocalNet [38] (d) Ours (e) GT
IV-C2 Effectiveness of HAFM
To assess the role of HAFM, we design four variants: HAFM-S (spatial modulation only), HAFM-F (frequency modulation only), HAFM-SSA (standard spatial attention replacing dark channel-guided attention), and HAFM-SMLP (spatial MLP replacing frequency modulation). The results are presented in Table III.
According to Table III, using only spatial (HAFM-S) or only frequency modulation (HAFM-F) results in PSNR drops of 1.06 dB and 1.00 dB, respectively, compared to the full HAFM. This underscores the importance of combining both spatial and frequency modulation for optimal performance. Replacing dark channel-guided attention with standard spatial attention (HAFM-SSA) leads to a 0.40 dB drop in PSNR and a 0.001 decrease in SSIM, highlighting that the haze-specific spatial modulation offers a stronger prior to frequency modulation. Substituting the frequency-domain MLP with a spatial MLP (HAFM-SMLP) increases FLOPs by 48.1%, yet PSNR remains 0.27 dB lower than the full HAFM, showing that frequency-domain processing not only boosts computational efficiency but also enhances global feature modeling.
IV-C3 Effectiveness of MGAM
To assess the contribution of each MGAM component, we conduct an ablation study with four variants: MGAM-$3\times 3$ (only the $3\times 3$ gated branch), MGAM-$5\times 5$ (only the $5\times 5$ gated branch), MGAM-Nogate (both branches without gating), and MGAM-Noskip (removing the skip connection). The results are reported in Table IV.
TABLE V: Ablation study for key components in PCGB on SOTS-Indoor.
@@ IV-C3 Effectiveness of MGAM
TABLE VI: Ablation study for key components in PCGB on SOTS-Outdoor.
@@
IV-C4 Effectiveness of PCGB
To validate the rationale behind PCGB, we test four variants: (1) PCGB-SSA, which employs standard spatial attention, identical to HAFM-SSA; (2) PCGB-NoFR, where the spatial attention map is generated from the initial dark channel features without feedback refinement (3) PCGB-DAF, which directly adds feedback correction to the initial dark channel features, bypassing SKFusion [14]; (4) PCGB-PFF, which uses the initial prior only in the first DGFDBlock, while subsequent blocks fuse the guidance and feedback from the previous DGFDBlock via SKFusion. Results on the SOTS-Indoor and SOTS-Outdoor datasets are exhibited in Table V and Table VI, respectively.
@@ IV-C4 Effectiveness of PCGB
(a) Hazy image (b) GT (c) Dark channel (d) First HAFM Attnmap (e) Last HAFM Attnmap
@@
Indoor results. Table V shows that the dark channel prior remains a reliable guidance cue in indoor scenes. Even without feedback refinement, PCGB-NoFR outperforms PCGB-SSA by 0.12 dB in PSNR, demonstrating the prior’s effectiveness in haze localization and degradation estimation. Adding feedback correction (PCGB-DAF) improves PSNR by 0.12 dB, and using SKFusion in the full PCGB increases both PSNR by 0.16 dB and SSIM by 0.001, demonstrating better adaptability to complex structures. However, PCGB-PFF, which discards the prior after the first block, results in a 0.71 dB PSNR drop, highlighting the importance of consistently using the initial prior for stable and accurate guidance.
Outdoor results. In outdoor environments, the dark channel prior becomes less reliable due to factors like sky and bright regions. As shown in Table VI, PCGB-NoFR performs 0.28 dB worse than PCGB-SSA, highlighting the limitations of the prior alone. Introducing feedback correction (PCGB-DAF) improves performance, yielding gains of 0.42 dB over PCGB-NoFR and 0.14 dB over PCGB-SSA. PCGB-PFF, which progressively fuses the prior with feedback, improves over NoFR by 0.29 dB and matches PCGB-SSA, proving the benefit of self-correction on noisy priors in outdoor scenes. However, PCGB-DAF still outperforms PCGB-PFF by 0.13 dB, emphasizing the importance of retaining the initial prior to avoid accumulating errors. The full PCGB achieves the best, surpassing PCGB-DAF by 0.56 dB, validating SKFusion’s ability to handle complex outdoor haze patterns.
IV-C5 Visual Analysis of PCGB
To further evaluate the corrective effect of PCGB, we visualize the dark channel-guided spatial attention maps from the first and last HAFM modules of DGFDNet. The visualization covers three representative scenes from the SOTS-Indoor, SOTS-Outdoor, and NH-HAZE datasets, as shown in Fig. 8.
In the indoor scene, the initial attention map follows the raw dark channel distribution, causing widespread activation. After feedback refinement, attention is focused on heavily degraded areas, improving haze localization and spatial modulation. In the outdoor scene, the limitations of dark channel priors are clear. In bright regions like the sky, the initial map overemphasizes these areas, neglecting distant buildings and roads. Feedback refinement shifts attention to the other haze regions, reducing over-processing of the sky.
In the NH-HAZE example, the initial attention map incorrectly activates on irrelevant structures, like the swing’s mesh, causing false positives. It also focuses on dense haze areas while missing subtly degraded regions. After refinement, the final map better aligns with the haze distribution, suppressing false activations and improving sensitivity to neglected regions. These results demonstrate the robustness of PCGB in handling complex, spatially diverse degradation patterns.
V Conclusion
@@
Acknowledgments
This work was partially supported by National Natural Science Foundation of China (Nos.62471317), Natural Science Foundation of Shenzhen (No. JCYJ20240813141331042), and Guangdong Provincial Key Laboratory (Grant 2023B1212060076).

### v3

@@ 4.2.2 Qualitative Comparisons
In Figure 7, two dense haze scenes are compared. In the first, our method achieves slightly lower PSNR than Dehamer [16], but surpasses it by 0.183 in SSIM, offering clearer textures and reduced color distortion. In the second scene, with severe haze and significant detail loss, our method restores the overall structure and preserves more details than Dehamer [16] and MB-TaylorFormer-B [29].
@@
Figure 8 compares two challenging non-homogeneous haze scenes. Our method achieves the highest scores across all metrics (PSNR and SSIM) in both scenes, demonstrating comprehensive superiority. In the first scene, it minimizes color distortion and avoids heavy artifacts from other methods. In the second scene, it removes haze most effectively, revealing clearer details, though it inadvertently eliminates some snow traces.
@@
Overall, our method attains the highest SSIM across all real-world cases, indicating strong global structural recovery even in the presence of severe haze-induced detail degradation.
4.3 Ablation Studies
Table 2: Ablation study for three core modules.
@@ 4.3 Ablation Studies
4.3.1 Effectiveness of Core Modules
To thoroughly evaluate the effectiveness and complementarity of the three core modules, we design four variants: MGAM, HAFM (without feedback correction), HAFM + PCGB (with full feedback-guided correction), and MGAM + HAFM. Table 2 summarizes the performance of each variant.
As shown in Table 2, MGAM improves PSNR by 0.75 dB over HAFM while maintaining the same SSIM of 0.994, highlighting its advantage in enhancing local details. Introducing PCGB into HAFM yields an additional 0.70 dB gain, emphasizing its role in enhancing global degradation modeling via adaptive prior correction. Combining MGAM and HAFM results in a 1.80 dB PSNR improvement over MGAM alone and 2.55 dB over HAFM, benefiting from the synergy between multi-scale spatial feature extraction and dual-domain modulation. Further adding PCGB achieves the best overall performance, confirming that the three modules complement each other and jointly contribute to both local and global restoration with minimal additional computational cost.
4.3.2 Effectiveness of HAFM
To assess the role of HAFM, we design four variants: HAFM-S (spatial modulation only), HAFM-F (frequency modulation only), HAFM-SSA (standard spatial attention replacing dark channel-guided attention), and HAFM-SMLP (spatial MLP replacing frequency modulation). The results are presented in Table 3.
Table 3: Ablation study for key components in HAFM.
@@ 4.3.2 Effectiveness of HAFM
According to Table 3, using only spatial (HAFM-S) or only frequency modulation (HAFM-F) results in PSNR drops of 1.06 dB and 1.00 dB, respectively, compared to the full HAFM. This underscores the importance of combining both spatial and frequency modulation for optimal performance. Replacing dark channel-guided attention with standard spatial attention (HAFM-SSA) leads to a 0.40 dB drop in PSNR and a 0.001 decrease in SSIM, highlighting that the haze-specific spatial modulation offers a stronger prior to frequency modulation. Substituting the frequency-domain MLP with a spatial MLP (HAFM-SMLP) increases FLOPs by 48.1%, yet PSNR remains 0.27 dB lower than the full HAFM, showing that frequency-domain processing not only boosts computational efficiency but also enhances global feature modeling.
4.3.3 Effectiveness of MGAM
To assess the contribution of each MGAM component, we conduct an ablation study with four variants: MGAM-$3\times 3$ (only the $3\times 3$ gated branch), MGAM-$5\times 5$ (only the $5\times 5$ gated branch), MGAM-Nogate (both branches without gating), and MGAM-Noskip (removing the skip connection). The results are reported in Table 4.
Table 4: Ablation study for key components in MGAM.
@@ 4.3.3 Effectiveness of MGAM
4.3.4 Effectiveness of PCGB
To validate the rationale behind PCGB, we test four variants: (1) PCGB-SSA, which employs standard spatial attention, identical to HAFM-SSA; (2) PCGB-NoFR, where the spatial attention map is generated from the initial dark channel features without feedback refinement (3) PCGB-DAF, which directly adds feedback correction to the initial dark channel features, bypassing SKFusion [31]; (4) PCGB-PFF, which uses the initial prior only in the first DGFDBlock, while subsequent blocks fuse the guidance and feedback from the previous DGFDBlock via SKFusion. Results on the SOTS-Indoor and SOTS-Outdoor datasets are exhibited in Table 5 and Table 6, respectively.
Table 5: Ablation study for key components in PCGB on SOTS-Indoor.
@@ 4.3.4 Effectiveness of PCGB
Table 6: Ablation study for key components in PCGB on SOTS-Outdoor.
@@
Indoor results. Table 5 shows that the dark channel prior remains a reliable guidance cue in indoor scenes. Even without feedback refinement, PCGB-NoFR outperforms PCGB-SSA by 0.12 dB in PSNR, demonstrating the prior’s effectiveness in haze localization and degradation estimation. Adding feedback correction (PCGB-DAF) improves PSNR by 0.12 dB, and using SKFusion in the full PCGB increases both PSNR by 0.16 dB and SSIM by 0.001, demonstrating better adaptability to complex structures. However, PCGB-PFF, which discards the prior after the first block, results in a 0.71 dB PSNR drop, highlighting the importance of consistently using the initial prior for stable and accurate guidance.
Outdoor results. In outdoor environments, the dark channel prior becomes less reliable due to factors like sky and bright regions. As shown in Table 6, PCGB-NoFR performs 0.28 dB worse than PCGB-SSA, highlighting the limitations of the prior alone. Introducing feedback correction (PCGB-DAF) improves performance, yielding gains of 0.42 dB over PCGB-NoFR and 0.14 dB over PCGB-SSA. PCGB-PFF, which progressively fuses the prior with feedback, improves over NoFR by 0.29 dB and matches PCGB-SSA, proving the benefit of self-correction on noisy priors in outdoor scenes. However, PCGB-DAF still outperforms PCGB-PFF by 0.13 dB, emphasizing the importance of retaining the initial prior to avoid accumulating errors. The full PCGB achieves the best, surpassing PCGB-DAF by 0.56 dB, validating SKFusion’s ability to handle complex outdoor haze patterns.
@@
(a) Hazy image (b) GT (c) Dark channel (d) First Attn (e) Last Attn
@@
4.3.5 Visual Analysis of PCGB
To further evaluate the corrective effect of PCGB, we visualize the dark channel-guided spatial attention maps from the first and last HAFM modules of DGFDNet. The visualization covers three representative scenes from the SOTS-Indoor, SOTS-Outdoor, and NH-HAZE datasets, as shown in Figure 9.
In the indoor scene, the initial attention map closely reflects the raw dark channel distribution, resulting in widespread activation across the image. After feedback refinement, the attention becomes more concentrated on heavily degraded areas, improving both the accuracy of haze localization and the effectiveness of spatial modulation.
In contrast, the outdoor scene reveals the limitations of dark channel priors. Due to low dark channel values in bright areas such as the sky, the initial attention map tends to overemphasize these regions while neglecting genuinely hazy areas like distant buildings and road surfaces. Through feedback refinement, PCGB progressively redirects attention away from the sky and toward the actual degradation, narrowing the response gap and mitigating over-processing of the sky.
In the NH-HAZE example, the initial attention map mistakenly activates on irrelevant structures such as the swing’s mesh, resulting in false positives. Moreover, it primarily focuses on dense haze areas while neglecting mildly degraded regions, leading to insufficient coverage. After iterative refinement, the final attention map more accurately aligns with the actual haze distribution. It effectively suppresses false activations and enhances sensitivity to previously neglected regions. These results confirm the robustness of PCGB in handling spatially diverse and challenging degradation patterns.
5 Conclusion
@@
Acknowledgment
This work was partially supported by National Natural Science Foundation of China (No.62471317), Natural Science Foundation of Shenzhen (No. JCYJ20240813141331042), and Guangdong Provincial Key Laboratory (Grant 2023B1212060076).